# Joint semi-supervised and contrastive learning enables zero-shot domain-adaptation and multi-domain segmentation

### v1

@@ IV-A Joint learning for unsupervised domain adaptation across imaging devices
UpperBound 61.30$\pm$ 4.34 N/A 9.61$\pm$ 1.42 N/A
Baseline 39.23$\pm$ 8.18 67.14$\pm$ 8.24 16.47$\pm$ 2.15 6.00$\pm$ 2.32
[…]
SimSiam 41.14$\pm$ 7.86 67.40$\pm$ 8.63 15.98$\pm$ 1.89 6.03$\pm$ 2.44
[…]
SegCLR($P_{\mathrm{s}}$,$C_{\mathrm{pool}}$) 39.39$\pm$ 8.38 52.24$\pm$ 7.35 17.25$\pm$ 6.00 8.29$\pm$ 3.27
@@
Learning representations of $D^{\mathrm{t}}$ with SimCLR and SimSiam with subsequent finetuning on $D^{\mathrm{s}}$ shows a slight improvement over Baseline for $D^{t}$, as seen in Figure 5 by their ($\approx$2%) higher Dice and lower UVD. This confirms that these pretraining strategies apply also on our OCT data. The results for $D^{s}$ are comparable to Baseline, as the finetuning on this domain, i.e., initializing the Baseline with the target-learned representations, does not seem to add more information than existing from the source supervision.
Figure 5: Relative segmentation metrics for models using $D^{s}=D_{1}$ and ${D^{t}=D_{2}}$. The dashed line depicts the average Baseline result used as reference for the relative metrics.
Most configurations of SegCLR lead to results superior to other methods. In Table I, SegCLR($P_{\mathrm{a}}$,$C_{\mathrm{ch}}$) is seen to consistently produce the best or the second best segmentation results for both $D^{s}$ and $D^{t}$ for DA across imaging devices. In particular, this configuration is seen to preserve the source domain information most successfully, with the best metric in $D^{s}$ — notably even compared to the supervised Baseline. This shows that supervised information from the labeled domain is not forgotten (e.g., as a model capacity trade-off when learning from the large unlabeled domain) and is indeed even enhanced and enriched with the added contrastive strategy. Hence, SegCLR can be used for segmentation of both $D^{s}$ and $D^{t}$, replacing the conventional Baseline approaches for the labeled domains, while approaching UpperBound such that it obviates labeling efforts for scenarios with no training data is or can be available in a target domain.
IV-B Unsupervised domain adaptation across eye diseases
This section evaluates SegCLR for domain adaptation across eye diseases, i.e., when image appearances are similar but the content differs. To that end, an experimental setting with $D^{s}=D_{1}$ and $D^{t}=D_{3}$ is used. Similarly to the previous section, Figure 6 and Table II show that the Baseline model trained on $D^{s}$ performs substantially worse on $D^{t}$ compared to an UpperBound model trained on that target domain. Interestingly, in contrast to the previous section, SimCLR, i.e., a model with contrastively pretrained features, is worse than the Baseline for both $D^{s}$ and $D^{t}$ in this setting. This can be due to instabilities in the contrastive loss, which do not affect SegCLR thanks to the proposed joint training strategy.
Model Dice UVD
@@ IV-B Unsupervised domain adaptation across eye diseases
UpperBound 79.32$\pm$ 14.25 N/A 6.57$\pm$ 4.91 N/A
Baseline 64.77$\pm$ 11.71 67.14$\pm$ 8.24 9.38$\pm$ 7.06 6.00$\pm$ 2.32
[…]
SimSiam 65.71$\pm$ 12.23 66.90$\pm$ 7.81 9.21$\pm$ 6.87 5.95$\pm$ 2.33
[…]
SegCLR($P_{\mathrm{s}}$,$C_{\mathrm{pool}}$) 63.13$\pm$ 15.61 63.01$\pm$ 7.41 10.49$\pm$ 7.08 6.25$\pm$ 2.40
@@
SegCLR($P_{\mathrm{s}}$,$C_{\mathrm{ch}}$) 65.00$\pm$ 11.09 65.85$\pm$ 7.10 9.82$\pm$ 6.68 6.20$\pm$ 2.50
@@
Figure 6: Relative segmentation metrics for models using $D^{s}=D_{1}$ and ${D^{t}=D_{3}}$. The dashed line depicts the Baseline results used as reference for the relative metrics.
@@ IV-C Selecting the reference SegCLR implementation
SegCLR($P_{\mathrm{a}}$,$C_{\mathrm{ch}}$) 3.65 4.66 3.68 4.75 4.21
SegCLR($P_{\mathrm{s+a}}$,$C_{\mathrm{ch}}$) 3.61 4.88 3.67 4.67 4.24
SegCLR($P_{\mathrm{a}}$,$C_{\mathrm{pool}}$) 3.82 4.72 3.98 4.73 4.34
SegCLR($P_{\mathrm{s+a}}$,$C_{\mathrm{pool}}$) 3.52 4.81 4.24 4.63 4.38
SimSiam 6.17 4.37 4.89 4.71 4.91
SegCLR($P_{\mathrm{s}}$,$C_{\mathrm{ch}}$) 4.81 4.75 5.02 5.21 4.96
Baseline 6.51 4.47 5.11 4.68 5.05
SimCLR 6.21 4.64 5.95 5.71 5.59
SegCLR($P_{\mathrm{s}}$,$C_{\mathrm{pool}}$) 6.59 7.54 5.85 5.72 6.38
The rankings confirm the clear advantage of most SegCLR models relative to Baseline, SimCLR, and SimSiam. They also show that pretrained models do not help in such unspervised domain adaptation tasks, for which joint training with SegCLR appears to be the key (with the only exception for the use of the pairing strategy $P_{\mathrm{s}}$, as discussed above). Although the rankings differ slightly depending on whether the source or the target domain performance is prioritized, overall SegCLR($P_{\mathrm{a}}$,$C_{\mathrm{ch}}$) is seen to be the best performer. By not requiring slice sampling, this model also has a simpler implementation in practice, e.g., compared to SegCLR($P_{\mathrm{s+a}}$,$C_{\mathrm{ch}}$), which is a top contender. In addition, the proposed $C_{\mathrm{ch}}$ adds a mere 0.03% more parameters to the baseline model. Accordingly, we chose SegCLR($P_{\mathrm{a}}$,$C_{\mathrm{ch}}$) as the reference configuration for our proposed method, calling it simply SegCLR henceforth, to use in the subsequent experiments and ablations.
@@ IV-D Effect of amount of target data on SegCLR
The success of modern self-supervised learning frameworks has been largely attributed to the availability and access to large unlabeled datasets [12, 15, 14, 13]. Previous sections demonstrate the efficacy of SegCLR given such large datasets, from extensive clinical trials which are rarely available in practice. An essential question thus remains as to how such framework performs in scenarios with smaller sets of unlabeled data. This question is studied here as an ablation study by reducing the amount of unlabeled data. We evaluate this with the experimental setting for DA across imaging devices (Section IV-A).
@@
Figure 7 presents the results of SegCLR when trained with logarithmically reducing amount of unlabeled data compared to the earlier experiments. The experiments are taken down to the point of including no unlabaled data, i.e., the contrastive loss being computed only using the source domain data (jointly with its supervised loss). Results show that the source domain performance is not affected much by the amount of unlabeled data from a different domain, which is expected since the supervised source domain is the major source of information and the main driver for this task. Interestingly, however, regardless of the amount of unlabaled data, SegCLR performs roughly 1% in Dice on average compared to Baseline. This suggests that the proposed framework and contrastive strategy implicitly advances and improves standard UNet training and its generalization to test data from the same source domain, regardless of data existence from yet another domain.
@@
Results in Figure 7 for the target domain of interest shows an expected trend of deterioration with diminishing amount of target data. Nevertheless, surprisingly the reduction in performance is quite minor, and a reduction of only 5.25$\pm$7.33Dice and -0.99$\pm$2.13UVD is observed between no unlabeled target data and 100% unlabeled target data (of nearly 7 million B-scans). Strikingly, SegCLR without any unlabeled data (i.e., SegCLR zero-shot learning) performs still significantly superior to Baseline, by 13.84$\pm$6.22Dice and -3.85$\pm$1.64UVD. This again suggests that the introduced framework with the joint learning strategy inherently enables better generalization, even in supervised settings without requiring additional data and without adding much to the network complexity. We further study this surprising observation in the next section.
@@
Figure 7: Segmentation results of SegCLR for ablation experiment of unlabeled data size. The 95% confidence interval is shown as blue bands for SegCLR and as gray dotted lines for Baseline replicates. The average of Baseline replicates (black dashed line at 0) is the basis for relative comparisons, i.e., a rel. Dice above 0 or a rel. UVD below 0 indicates a superior performance by SegCLR.
IV-E Zero-shot domain adaptation on different domain combinations
@@
In this section, we assess SegCLR zero-shot DA capability across different experimental settings. To that end, we consider all possible combinations of the datasets $D_{1}$, $D_{2}$, and $D_{3}$ as $D^{s}$ and $D^{t}$. Note that in this zero-shot setting, $D^{t}$ is exclusively reserved for evaluation, with its data (including unlabeled images) not been seen during the training phase.
@@
Figure 8 and Table IV demonstrate the superiority of SegCLR over Baseline (UNet) consistently in nearly all experimental dataset combinations. The only setting SegCLR is clearly inferior (for both Dice and UVD) is for training on $D^{s}=D_{3}$ and evaluating on $D^{t}=D_{1}$, for which the imaging device stays the same but the disease type (labeling info and label sets) mainly differ. Note that with the imaging device DA from $D^{s}\in\{D_{1},D_{2}\}$ to $D^{t}=D_{2}$ leads a relatively larger reduction in performance for Baseline, which cannot generalize across such large domain gap — a setting in which SegCLR appears to be relatively successful. Visual examples from different experimental combinations are shown in Figure 9.
Figure 8: Results for zero-shot domain adaptation models trained on different source domains $D^{s}$ shown in the x axes, with $D^{\mathrm{All}}=D_{1}\cup D_{2}\cup D_{3}$. Each graph shows the evaluation results on each target domains $D^{t}$. The metrics are relative to the average of the Baseline models trained on the same $D^{s}$. The distributions of Baselines are also provided to illustrate the variation in Baseline repetitions. A positive Dice value for training on $D_{i}$ and evaluating on $D_{j}$ means that a zero-shot CL with SegCLR from domain $i$ applied to domain $j$ is superior to conventional supervised-only training. This indicates that the domain gap between $i$ and $j$ can be bridged by representation learning with SegCLR. Note that, although not directly observable in these normalized plots, the absolute Baseline performance for $i=j$ is much higher than otherwise, as expected. These results for $i=j$ show that SegCLR is superior to traditional supervised learning Baseline, even when applied on the trained domain itself, i.e., without any domain shift. SegCLR can thus augment and replace conventional UNet for better generalizability.
@@
Overall, SegCLR is seen to outperform Baseline across different combinations of source and target datasets. Additionally, SegCLR surpasses Baseline even when evaluated within the same source domain. This suggests that the benefits of SegCLR extend beyond DA alone, proving valuable even for conventional supervised settings with training and deployment within the same domain, i.e., as a plug-and-play replacement for UNet.
@@
Collectively, these results affirm the prior hypothesis that SegCLR exhibits robust performance in realistic scenarios where deployment on entirely unseen domains is anticipated, particularly when the domain gap arises from changes image apperance, e.g., across different imaging devices.
@@
Figure 9: Qualitative assessment of SegCLR and Baseline, trained on different domains (rows) for example images from each domain (columns). Note that $D_{3}$ only has two of the four labels. These visually examplify that ($i$) training and evaluating on the same domain works well with both Baseline and SegCLR, ($ii$) SegCLR outperforms Baseline when training and evaluation domains differ, i.e., in UDA settings, ($iii$) multi-domain training $D_{\mathrm{All}}$ performs better than single-domain training, for both Baseline and SegCLR, with minor superiority of SegCLR (see bottom-right for $D_{3}$).

### v3

@@ IV-A Joint learning for unsupervised domain adaptation across imaging devices
For state-of-the-art comparison two additional approaches that does not utilize self-supervised learning are included. CycleGAN [30] was adapted to our UNet for OCT slices and it was trained to translate images of $D^{t}$ to the domain of $D^{s}$. On such translated images, we then applied the Baseline UNet trained on $D^{s}$. We qualitatively confirmed the convergence of training that yields meaningful translated images. In Figure 5 the CycleGAN results for $D^{t}$ are seen to be slightly inferior to Baseline, contrary to the observation reported in [30]. This is probably due to their baseline results having a near-zero Dice, while our Baseline being a relatively much superior contender (see Table I). We also include singular value decomposition-based noise adaptation (SVDNA) [31], which was published concurrently with our preliminary results of SegCLR [25]. SVDNA is a style transfer technique that utilizes difference in noise structure between domains. We used its public-domain implementation [32], by adapting its inherent parameter $k$, which denotes the threshold of singular values used, as $k\in[40,100]$ to accommodate for our different image size. The results for SVDNA in Figure 5 are superior to Baseline in the target domain. Note that such methods for image/style transfer to a target domain are not intended for applying on the source domain, which is corroborated by the SVDNA results in Figure 5 having similar UVD and lower Dice compared to Baseline for $D^{s}$. The CycleGAN model is outright not applicable for the source domain.
Standard self-supervised learning methods learn the representation of $D^{\mathrm{t}}$. Finetuning this subsequently on $D^{\mathrm{s}}$, SimCLR and SimSiam achieve a slight improvement over Baseline for $D^{t}$, as seen in Figure 5 by their ($\approx$2%) higher Dice and lower UVD. This confirms that these pretraining strategies apply also on our OCT data. The results for $D^{s}$ are comparable to Baseline, as the finetuning on this domain, i.e., initializing the Baseline with the target-learned representations, does not seem to add more information than existing from the source supervision. Note that for SimCLR and SimSiam we use their standard implementations including their original sample pairing strategy (corresponding to $P_{\mathrm{a}}$ in our terminology) and contrastive-head aggregation approach (equivalent to $C_{\mathrm{pool}}$).
Most configurations of SegCLR lead to results superior to Baseline. In Table I, SegCLR($P_{\mathrm{a}}$,$C_{\mathrm{ch}}$) is seen to produce the best or the second best segmentation results in most cases for both $D^{s}$ and $D^{t}$ for DA across imaging devices. In particular, this configuration is seen to preserve the source domain information most successfully, with the best metric in $D^{s}$ — notably even compared to the supervised Baseline. This shows that supervised information from the labeled domain is not forgotten (e.g., as a model capacity trade-off when learning from the large unlabeled domain) and is indeed even enhanced and enriched with the added contrastive strategy. This is a feature that previous UDA methods such as CycleGAN and SVDNA do not have, as they aim at application on a specific target domain. Hence, SegCLR can be used for segmentation of both $D^{s}$ and $D^{t}$, replacing the conventional Baseline approaches for the labeled domains, while approaching UpperBound such that it obviates labeling efforts for scenarios with no training data is or can be available in a target domain.
@@
UNet (UpperBound) 61.30$\pm$ 4.34 N/A 961$\pm$ 142 N/A
UNet (Baseline) 39.23$\pm$ 8.18 67.14$\pm$ 8.24 1647$\pm$ 215 600$\pm$ 232
@@
SimSiam 41.14$\pm$ 7.86 67.40$\pm$ 8.63 1598$\pm$ 189 603$\pm$ 244
@@
SegCLR($P_{\mathrm{s}}$,$C_{\mathrm{pool}}$) 39.39$\pm$ 8.38 52.24$\pm$ 7.35 1725$\pm$ 600 829$\pm$ 327
@@
SegCLR($P_{\mathrm{s}}$,$C_{\mathrm{ch}}$) 50.60$\pm$ 8.98 66.38$\pm$ 7.06 1325$\pm$ 370 613$\pm$ 241
@@
Figure 5: Relative segmentation metrics for cross-device domain adaptation, i.e., using $D^{s}=D_{1}$ and ${D^{t}=D_{2}}$. The dashed line depicts the average Baseline result used as reference for the relative metrics. SVDNA is included in brackets for $D^{s}$, as in practice Baseline model would be used instead.
IV-B Unsupervised domain adaptation across eye diseases
This section evaluates SegCLR for domain adaptation across eye diseases, i.e., when image appearances are similar but the content differs. To that end, an experimental setting with $D^{s}=D_{1}$ and $D^{t}=D_{3}$ is used. Similarly to the previous section, Figure 6 and Table II show that the Baseline model trained on $D^{s}$ performs substantially worse on $D^{t}$ compared to an UpperBound model trained on that target domain. As in the previous subsection, CycleGAN results are similar to Baseline. In contrast, SVDNA performs only slightly better than Baseline in the target domain, while it was clearly superior in the previous section. This is in line with the emphasis in [31] that SVDNA as a noise adaptation method is successful mainly when the styles differ, e.g., between different devices as in the previous subsection, whereas it seems not effective in scenarios with similar appearances but different content between domains as in this subsection. Interestingly, in contrast to the previous section, SimCLR, i.e., a model with contrastively pretrained features, is worse than Baseline for both $D^{s}$ and $D^{t}$ in this setting. This can be due to instabilities in the contrastive loss, which do not affect SegCLR thanks to the proposed joint training strategy.
@@ IV-B Unsupervised domain adaptation across eye diseases
Figure 6: Relative segmentation metrics for cross-disease domain adaptation, i.e., using $D^{s}=D_{1}$ and ${D^{t}=D_{3}}$. The dashed line depicts the Baseline results used as reference for the relative metrics. SVDNA is included in brackets for $D^{s}$, as in practice Baseline model would be used instead.
Model Dice [%] UVD [fL]
[…]
UNet (UpperBound) 79.32$\pm$ 14.25 N/A 657$\pm$ 491 N/A
UNet (Baseline) 64.77$\pm$ 11.71 67.14$\pm$ 8.24 938$\pm$ 706 600$\pm$ 232
[…]
SVDNA 67.39$\pm$ 12.14 66.56$\pm$ 8.15 902$\pm$ 698 608$\pm$ 239
[…]
SimSiam 65.71$\pm$ 12.23 66.90$\pm$ 7.81 921$\pm$ 687 595$\pm$ 233
[…]
SegCLR($P_{\mathrm{s}}$,$C_{\mathrm{pool}}$) 63.13$\pm$ 15.61 63.01$\pm$ 7.41 1049$\pm$ 708 625$\pm$ 240
[…]
SegCLR($P_{\mathrm{s}}$,$C_{\mathrm{ch}}$) 65.00$\pm$ 11.09 65.85$\pm$ 7.10 982$\pm$ 668 620$\pm$ 250
[…]
@@ IV-C Selecting the reference SegCLR implementation
SegCLR($P_{\mathrm{a}}$,$C_{\mathrm{ch}}$) 4.08 5.56 4.57 5.60 5.01
SegCLR($P_{\mathrm{s+a}}$,$C_{\mathrm{ch}}$) 4.11 5.96 4.39 5.45 5.03
SegCLR($P_{\mathrm{a}}$,$C_{\mathrm{pool}}$) 4.28 5.65 4.98 5.54 5.19
SegCLR($P_{\mathrm{s+a}}$,$C_{\mathrm{pool}}$) 4.00 5.97 5.29 5.50 5.32
SVDNA 4.59 5.41 5.89 5.83 5.55
Baseline 7.90 5.41 5.68 5.83 6.00
SimSiam 7.34 5.17 6.55 5.67 6.09
SegCLR($P_{\mathrm{s}}$,$C_{\mathrm{ch}}$) 5.66 5.91 6.52 6.42 6.20
CycleGAN 8.66 5.41 6.46 5.83 6.37
SimCLR 7.44 5.69 8.26 7.26 7.21
SegCLR($P_{\mathrm{s}}$,$C_{\mathrm{pool}}$) 7.93 9.86 7.40 7.08 8.04
The rankings confirm the clear advantage of most SegCLR models relative to Baseline, SimCLR, and SimSiam. They also show that pretrained models do not help in such unsupervised domain adaptation tasks, for which joint training with SegCLR appears to be the key (with the only exception for the use of the pairing strategy $P_{\mathrm{s}}$, as discussed above). For ranking of CycleGAN and SVDNA on $D^{s}$, the Baseline results are used in their stead, not to disadvantage these methods as this would also be the practical deployment scenario. As mentioned above, CycleGAN is not applicable to $D^{s}$. While SVDNA is applicable, it was designed for $D^{t}$ only, as confirmed by the results above in Figure 5 and Figure 6, where its performance is comparable to or worse than Baseline. While CycleGAN ranks (6.37) slightly worse than Baseline (6.00), the improvement of SVDNA on $D^{t}$ makes it rank slightly higher overall (5.55). Despite being a close contender, SVDNA ranks behind most SegCLR configurations for UDA, i.e., in $D^{t}$ rankings including cross-device ($D^{t}=D_{2}$) setting, for which SVDNA was originally designed for. While SVDNA can merely mirror Baseline results for $D^{s}$ evaluations, SegCLR can leverage target domain info back for the source domain, facilitating its success for $D^{s}$ in the cross-disease ($D^{t}=D_{3}$) setting. Overall, most SegCLR configurations perform superior to SimSiam, SimCLR, CycleGAN, and SVDNA.
Although the rankings differ slightly depending on whether the source or the target domain performance is prioritized, overall SegCLR($P_{\mathrm{a}}$,$C_{\mathrm{ch}}$) is seen to be the best performer. By not requiring slice sampling, this model also has a simpler implementation in practice, e.g., compared to SegCLR($P_{\mathrm{s+a}}$,$C_{\mathrm{ch}}$), which is a top contender. In addition, the proposed $C_{\mathrm{ch}}$ adds a mere 0.03% more parameters to the baseline model. Accordingly, we chose SegCLR($P_{\mathrm{a}}$,$C_{\mathrm{ch}}$) as the reference configuration for our proposed method, calling it simply SegCLR henceforth, to use in the subsequent experiments and ablations.
@@ IV-D Effect of amount of target data on SegCLR
The success of modern self-supervised learning frameworks has been largely attributed to the availability and access to large unlabeled datasets [13, 16, 15, 14]. Previous sections demonstrate the efficacy of SegCLR given such large datasets, from extensive clinical trials which are rarely available in practice. An essential question thus remains as to how such framework performs in scenarios with smaller sets of unlabeled data. This question is studied here as an ablation study by reducing the amount of unlabeled data on the target domain $D^{t}$. We evaluate this with the experimental setting for DA across imaging devices (Section IV-A), i.e., $D^{s}=D_{1}$ and $D^{t}=D_{2}$.
@@
Figure 7 presents the results of SegCLR when trained with logarithmically reducing amount of unlabeled data compared to the earlier experiments. The experiments are taken down to the point of including no unlabeled data, i.e., the contrastive loss being computed only using the source domain data (jointly with its supervised loss). Results show that the source domain performance is not affected much by the amount of unlabeled data from a different domain, which is expected since the supervised source domain is the major source of information and the main driver for this task. Interestingly, however, regardless of the amount of unlabeled data, SegCLR performs roughly 1% in Dice on average compared to Baseline. This suggests that the proposed framework and contrastive strategy implicitly advances and improves standard UNet training and its generalization to test data from the same source domain, regardless of data existence from yet another domain.
@@
Results in Figure 7 for the target domain of interest shows an expected trend of deterioration with diminishing amount of target data. Nevertheless, surprisingly the reduction in performance is quite minor, and a reduction of only 5.25$\pm$7.33Dice and -0.99$\pm$2.13UVD is observed between no unlabeled target data and 100% unlabeled target data (of nearly 7 million B-scans). Strikingly, SegCLR without any unlabeled data (i.e., SegCLR for domain generalization) performs still significantly superior to Baseline, by 13.84$\pm$6.22Dice and -3.85$\pm$1.64UVD. This again suggests that the introduced framework with the joint learning strategy inherently enables better generalization, even in supervised settings without requiring additional data and without adding much to the network complexity. This is indeed a remarkable feature of our proposed method in contrast to previous UDA methods (e.g., CycleGAN and SVDNA), which by design require unlabeled data from the target domain. We further study this surprising observation in the next section.
@@
Figure 7: Segmentation results of SegCLR for ablation experiment, where unlabeled data on the target domain $D^{t}$ is systematically reduced. The 95% confidence interval is shown as blue bands for SegCLR and as gray dotted lines for Baseline replicates. The average of Baseline replicates (black dashed line at 0) is the basis for relative comparisons, i.e., a rel. Dice above 0 or a rel. UVD below 0 indicates a superior performance by SegCLR. Note the different y-axis scales of the source and target domain plots.
IV-E Domain generalization on different domain combinations
@@
In this section, we assess SegCLR domain generalization capability across different experimental settings. To that end, we consider all possible combinations of the datasets $D_{1}$, $D_{2}$, and $D_{3}$ as $D^{s}$ and $D^{t}$. Additionally, we perform evaluations on a combined dataset $D_{\mathrm{All}}=D_{1}\cup D_{2}\cup D_{3}$ as $D^{s}$, which is motivated and detailed further in the next section. Note that in this domain generalization setting, $D^{t}$ is exclusively reserved for evaluation, with its data (including unlabeled images) not been seen during the training phase.
@@
Figure 8 and Table IV demonstrate the superiority of SegCLR over UNet consistently in nearly all experimental dataset combinations. The only setting SegCLR is clearly inferior (for both Dice and UVD) is for training on $D^{s}=D_{3}$ and evaluating on $D^{t}=D_{1}$, for which the imaging device stays the same but the disease type (labeling info and label sets) mainly differ. Note that with the imaging device DA from $D^{s}\in\{D_{1},D_{2}\}$ to $D^{t}=D_{2}$ leads a relatively larger reduction in performance for UNet, which cannot generalize across such large domain gap — a setting in which SegCLR appears to be relatively successful. Visual examples from different experimental combinations are shown in Figure 9.
Figure 8: Results for domain generalization models trained on different source domains $D^{s}$ shown in the x axes, with $D_{\mathrm{All}}=D_{1}\cup D_{2}\cup D_{3}$. Each row of plots shows the evaluation results on a target domain $D^{t}$. In the supervised learning setting when the evaluation domain matches the training domain (indicated by bold x-labels), UNet results correspond to UpperBound per our previous naming. In other settings, these UNet results are the Baseline, i.e., application without any specific domain adaptation effort. Although the margin for improvement becomes small with both models performing well in this supervised setting, SegCLR can surprisingly still outperform UNet in both Baseline and UpperBound settings. SegCLR can thus augment and replace conventional UNet for better generalizability.
@@
Overall, SegCLR is seen to outperform the UNet baseline across different combinations of source and target datasets. Additionally, SegCLR surpasses UNet even when evaluated within the same source domain, the scenario denoted as UNet (UpperBound) in the previous sections. This suggests that the benefits of SegCLR extend beyond DA alone, proving valuable even for conventional supervised settings with training and deployment within the same domain, i.e., as a plug-and-play replacement for UNet.
@@
Collectively, these results affirm the prior hypothesis that SegCLR exhibits robust performance in realistic scenarios where deployment on entirely unseen domains is anticipated, particularly when the domain gap arises from changes image appearance, e.g., across different imaging devices.
@@
Figure 9: Qualitative assessment of SegCLR and UNet baseline, trained on different domains (rows) for example images from each domain (columns). Note that $D_{3}$ only has two of the four labels. These visually exemplify that ($i$) training and evaluating on the same domain works well with both UNet and SegCLR, ($ii$) SegCLR outperforms UNet when training and evaluation domains differ, i.e., in UDA settings, ($iii$) multi-domain training $D_{\mathrm{All}}$ performs better than single-domain training, for both UNet and SegCLR, with minor superiority of SegCLR (see bottom-right for $D_{3}$).